OpenStudy (anonymous):

The solution to the radical function is x = 6.Answer True False

OpenStudy (anonymous):



OpenStudy (anonymous):

@ghazi

OpenStudy (anonymous):

Just switch x for 6 and see if you obtain the result

OpenStudy (anonymous):

I got false.

OpenStudy (ghazi):

\[\sqrt{\frac{ 1 }{ 2 }x+1}=2\] now square both sides and you dont need to put x=6 , thats the question

OpenStudy (anonymous):

I did get false by the way...

OpenStudy (ghazi):

\[\frac{ 1 }{ 2 }x+1=4\] then \[\frac{ 1 }{ 2 }x=4-1=3\] and finally x=3*2=6 so its true

OpenStudy (anonymous):

Fault

OpenStudy (anonymous):

my fault gosh how I miss that...

OpenStudy (ghazi):

see here is the total solution you should have waited